OpenStudy (anonymous):

find the vertex, line of symmetry and range, -x^2-4x-1

OpenStudy (anonymous):

\[\Huge \frac{-b}{2a}\]

OpenStudy (anonymous):

I need someone to walk me thru

OpenStudy (anonymous):

I think vertex(-2,3) line of symmetry x=-2

OpenStudy (anonymous):

I didn't get that

OpenStudy (anonymous):

let y=-x^2-4x-1 y=-((x+2)^2-3) so vertex is (-2,3) and it'll be symmetrical to line x=-2

OpenStudy (anonymous):

and range is (-infnity,3)

OpenStudy (anonymous):

I don't see where you get -3 in your y=((x+2)^2-3)

OpenStudy (anonymous):

take - common in expression and 3 and -3 in expression so y=-(x^2+4x+1+3-3) y=-(x^2+4x+4-3), so y=-((x+2)^2-3)

OpenStudy (anonymous):

Where does +3 and -3 come from????

OpenStudy (anonymous):

in any algebric expression ,if you add "a" and subtract "a" then it does not change the expression

OpenStudy (anonymous):

I appreciate your help but I don't get it. Thank you for your time.